# Implementation Study of Cost-Effective Verification for Pietrzak’s Verifiable Delay Function in Ethereum Smart Contracts

### v1

@@ 2.1.1 Definition and Fundamental Properties
2.2 Feasibility of Applying VDFs to Smart Contracts
In both Wesolowski [2] and Pietrzak [1], the prover claims $y=x^{2^{T}}\mod N$ where N is a RSA modulus, but the way to verify it is different. In Wesolowski, the verifier gives the prover a random and large prime challenge value c, and the prover gives the verifier a proof $\pi$ which is $x^{{2^{T}}/c}$. Here, the prover does not compute $\pi$ anew but rather responds by taking the intermediate value of the process of computing $y=x^{2^{T}}\mod N$ from memory. The verifier then checks if $\pi\in\mathbb{Z}_{N}^{*}$ and accepts if $y={\pi^{c}}{x^{{2^{T}}\mod c}}$. To make this scheme non-interactive in a blockchain environment, the Fiat-Shamir heuristic is utilized: the prover computes $c$ directly as hash-to-prime(x, y) and passes it along with $\pi$ to the verifier in the first message. To ensure the difficulty in creating a verifiable proof, it is computationally assumed that calculating $z^{c}=y\mod N$ is challenging for any $z\neq 1$. Although this operation is time-independent with respect to $T$, it requires considerable resources and presents significant implementation challenges. These include incorporating an on-chain hash-to-prime function capable of handling relatively large data sizes, such as 2048 bits.
On the other hand, Pietrzak VDF does not require techniques like hash-to-prime and makes no computational assumptions. In the Pietrzak protocol, proof generation and verification are halving protocol that involves a loop $log(T)$ times, dividing $T$ by 2 until $T$ becomes 1. During each iteration of the loop in a proof generation, the prover computes ${v_{i}}=x^{2^{T/2}}$. The Fiat-Shamir heuristic is used to compute $r=Hash(x_{i},y_{i},v_{i})$ to make it non-interactive in a blockchain environment. And the prover and verifier calculate $(x_{i+1},y_{i+1})\leftarrow({{x_{i}}^{r}{v_{i}}},{{v_{i}}^{r}}{y_{i}})$ that satisfies $y_{x+i}=x_{i+1}^{2^{T/2}}$. Thus, a proof $\pi=\{v_{i}\}_{i\in[1,\log_{2}T]}$ of $\log_{2}T$ length is produced. When each element of the proof is over $2048$ bits, it is expected to be impractical for use in a blockchain due to the large size [3].

### v2

@@ 2.1.1 Definition and Fundamental Properties
2.2 Feasibility of On-Chain VDF Verification
Both Wesolowski [2] and Pietrzak [1] VDFs operate by transforming a value $x$ into $y=x^{2^{T}}\mod N$, where $N$ is an RSA modulus, although their verification methods differ. Scalability and implementation costs are significant challenges in blockchain systems. Employing the Fiat-Shamir heuristic, both methods facilitate non-interactive verification, suitable for blockchain environments. However, the computational intensity and storage demands for managing large numeric values pose substantial hurdles.
Wesolowski’s VDF maintains a constant proof length ($O(1)$), independent of the time delay parameter $T$, making it particularly advantageous where proof size is a concern. The implementation on blockchain platforms like Ethereum is particularly challenging due to the high cost associated with verifying or finding primes for large numbers, such as 2048 bits. This necessity significantly increases the implementation cost, making full on-chain execution unfeasible.
Pietrzak’s VDF does not require operations like hash-to-prime, simplifying its on-chain implementation. The protocol involves a halving process, where $T$ is divided by two until it becomes one, reducing the complexity of computations with each iteration. However, the proof length and cost are influenced by the time delay variable $T$, and the proof length grows with $O(\log_{2}(T))$. This characteristic can result in verifications that exceed the size limit for feasible blockchain implementation. The research anticipated that the length of the proof will approximate 40 KB [3].